. The sound intensity levels for a machine shop and a quiet library are 95 dB and 39 dB, respectively. What is the difference between the intensity of the sound in the machine shop and that in the library?
Answer #1

Here,

as the dB level is given as

dB =10 * log(I1/Io)

I1 = Io * 10^(dB/10)

for the difference in inetensity

difference in inetensity = 1 *10^-12 * (10^(95/10) - 10^(39/10))

difference in inetensity = 3.162 *10^-3 W/m^2

the difference between the intensity of the sound in the machine shop and that in the library is 3.162 *10^-3 W/m^2